The scores on an national exam for 11th grade is are normally distributed with LaTeX: \mu=71 μ = 71 and LaTeX: \sigma=9 σ = 9 . Which of the following statements most precisely describes the probability that a randomly chosen 11th grader scored 90 or more on this test?
Answer #1


= 71

= 9

To findP(X90):

Z = (90-71)/9 = 2.1111

Table of Area Under Standard Normal Curve gives area = 0.4826

So,

P(X90) = 0.5 - 0.4826 = 0.0174

So,

Answer is:

0.0174